# The basis problem for modular forms for the Weil representation

### v1

@@ 4 Vector valued Eisenstein series
In this section we will study a relation between the Eisenstein series for genus $n=2$ and the Hecke operators for $n=1$. In the scalar valued case a similar result was shown in [G] and [], called the pullback formula. We generalize the pullback formula to the vector valued case, however using a different approach. In particular, we will prove that $\partial_{h}E_{m/2}^{(2)}\left(\left(\begin{smallmatrix}z&0\\ 0&z^{\prime}\end{smallmatrix}\right)\right)$ is the sum of the kernel functions for the Hecke operators from the previous section. When $h=0$ we get, as an additional term, the product of the genus $1$ Eisenstein series in $z$ and in $z^{\prime}$ (also cf. [St]).
@@
converges normally and thus defines a modular form (cf. [W2, Theorem 1]).
@@
Let $l\in\mathbb{Z}$, $A\in M_{l}$ and $B\in\mathrm{SL}_{2}(\mathbb{Z})$. Then $(AB)^{\prime}=B^{\prime}A^{\prime}$ and
@@
Let $\mathcal{D}$ be a discriminant form of even signature. And let $M\in\Gamma^{(2)}$ and $\epsilon=\operatorname{sgn}(\nu(M))$ with $\operatorname{sgn}(0)=-1$. Then
@@
Then $\psi$ defines a bijection between $\Gamma_{\infty}^{(1)}\backslash\Gamma^{(1)}\times\Gamma_{\infty}^{(1)}\backslash\Gamma^{(1)}$ and $M_{0}/\{\pm 1\}$ and we have
@@
The next result finally describes the relation between $E_{k}^{(2)}$ and the Hecke operators and is due to Stein (cf. [St, Theorem 5.3]). We include it together with its proof because we will generalize the argument to establish Theorem 4.7.
@@
where the second sum ranges over the appropriate $A$ and $B$ (which depend on $l$). It was shown in [I, section 3.1.1] that $\partial_{h}$ commutes with the slash operator of $u(A)$ and $d(B)$ on ×$\mathbb{H}\times\mathbb{H}\subset\mathbb{H}_{2}$, i.e. for a $f:\mathbb{H}_{2}\rightarrow\mathbb{C}$ we have
@@
5 Vector valued Siegel–Weil formula
In this section we will see that the genus theta series defined in section 2 is equal to the Eisenstein series defined in section 4. This is essentially an application of the Siegel–Weil formula shown by Weil in [W2]. Some missing cases were later completed by Kudla and Rallis in [KR]. The Siegel–Weil formula is formulated in an adelic setup that contains the classical case that we are interested in. We will state the Siegel–Weil formula and show how to deduce the equality of the genus theta series and the Eisenstein series from it. For this purpose we will first introduce this more general setup. We mostly follow [Li]. Most of the calculations in this section are known, but there seems to be no reference giving the result in the generality that we require.
@@ 5 Vector valued Siegel–Weil formula
$A$ =$=\left\{a(u)=\begin{pmatrix}u&0\\ 0&(u^{T})^{-1}\end{pmatrix}\bigm{|}u\in\operatorname{GL}_{n}\right\},$
$N$ =$=\left\{n(s)=\begin{pmatrix}I_{n}&s\\ 0&I_{n}\end{pmatrix}\bigm{|}s\in\mathrm{Sym}_{n}\right\}.$
Let $V$ be a quadratic space over $\mathbb{Q}$ of dimension $m=2k$ with non-degenerate symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) and $V(R)=V\otimes_{\mathbb{Q}}R$ for any ring $R\supset\mathbb{Q}$. Let $\operatorname{O}(V)$ be the corresponding orthogonal group scheme. Then $(\mathcal{G},\mathcal{H})=(\mathrm{Sp}_{2n},\operatorname{O}(V))$ is a reductive dual pair.
@@
given by mapping a basis element $e^{\underline{\mu}}$ of $\mathbb{C}[(L^{\prime}/L)^{n}]$ to $\varphi_{\underline{\mu}}=\bigotimes_{p<\infty}\varphi_{p}$, where $\varphi_{p}\in\mathscr{S}(V(\mathbb{Q}_{p})^{n})$ is the characteristic function of $\underline{\mu}+(L\otimes\mathbb{Z}_{p})^{n}$. We obtain (cf. [Zh, section 2.1.3])
@@
Theorem 5.4 (Siegel–Weil formula (cf. [W2] and [KR])).
@@
$\chi_{\infty}(\det U)^{-1}|\det U|_{\infty}^{-k}\cdot\int_{\mathcal{H}(\mathbb{Q})\backslash H(\mathbb{A})}\theta(g_{Z},h;\varphi)\mathrm{d}h$
@@
()\$\mathcal{H}(\mathbb{Q})\backslash H(\mathbb{A})/S\xrightarrow{\sim}G,\qquad h\mapsto h(L\otimes\hat{\mathbb{Z}})\cap V.$
@@
()\$\int_{\mathcal{H}(\mathbb{Q})\backslash H(\mathbb{A})}\theta(g_{Z},h;\varphi)\mathrm{d}h=\sum_{j}\int_{H(\mathbb{Q})\backslash H(\mathbb{Q})h_{j}S}\theta(g_{Z},h;\varphi)\mathrm{d}h.$
Substituting $h$ for $hh_{j}$ and applying $\mathcal{H}(\mathbb{Q})\cap h_{j}Sh_{j}^{-1}=\operatorname{Aut}(L_{j})$ we obtain
$\int_{H(\mathbb{Q})\backslash H(\mathbb{Q})h_{j}S}\theta\left(g_{Z},h;\varphi\right)\mathrm{d}h$ =()\$=\int_{\mathcal{H}(\mathbb{Q})\backslash H(\mathbb{Q})h_{j}Sh_{j}^{-1}}\theta\left(g_{Z},hh_{j};\varphi\right)\mathrm{d}h$
@@
Substituting $h$ for $h_{j}hh_{j}^{-1}$ the third integral becomes
@@
where in the last equation we used that the canonical homomorphism $S\rightarrow\operatorname{O}(\mathcal{D})$ is surjective (see [N, Corollary 1.9.6.]). Combining these computations we find
$\chi_{\infty}(\det U)^{-1}$ $|\det U|_{\infty}^{-k}\cdot\int_{\mathcal{H}(\mathbb{Q})\backslash H(\mathbb{A})}\theta(g_{Z},h;\varphi)\mathrm{d}h$
@@
$\mathrm{vol}(S)\cdot\mu(G)$ $=\mathrm{vol}(S)\cdot\sum_{j}\frac{1}{\#\operatorname{Aut}(L_{j})}=\mathrm{vol}(\mathcal{H}(\mathbb{Q})\backslash H(\mathbb{A}))=1$
@@
For details see for example [Ku]. We consider the finite part. Note that for $v\in\mathbb{C}[\mathcal{D}^{n}]$, $\underline{\beta}\in\mathcal{D}^{n}$ and $\underline{x}\in\underline{\beta}+L^{n}$ we have $\iota(v)(\underline{x})=\langle v,e^{\underline{\beta}}\rangle$. Furthermore, $\omega(g_{Z})$ acts trivially on $\iota(e^{\underline{\mu}})$ since $g_{Z}\in\mathcal{G}(\mathbb{R})$ and therefore
@@
6 Space of theta series
@@
Let $\mathcal{D}$ be a discriminant form of even signature and level $N$ and let $m$ be even with $m>p\text{-rank}(\mathcal{D})$ for all primes $p$. We set $G=I\!I_{m,0}(\mathcal{D})$ and $k=m/2+h$ with $h\geq 0$. By [N, Corollary 1.10.2] $G$ is non-empty. We define a linear map
@@
Let $V$ be a $\mathbb{C}$-vector space with scalar product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ), which is linear in the first variable and $(v_{i})_{i=1}^{n}\subset V$ an arbitrary finite family. Then $f:V\rightarrow V$ defined by
@@
and $g(x\cdot A)=f(g(x))$. From the first assertion we see that $\ker(g)=\ker(A)$ and since $A$ is self-adjoint $\mathbb{C}^{n}=\operatorname{im}(A)\oplus\ker(A)=\operatorname{im}(A)\oplus\ker(g)$. This implies that
@@
We remark that [St] derived the formula
$\int_{\Gamma\backslash\mathbb{H}}\langle f,E_{k}^{(2)}\left(\left(\begin{smallmatrix}z&0\\ 0&-\overline{z^{\prime}}\end{smallmatrix}\right),\overline{s}\right)\rangle y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}$ $=\frac{e(\operatorname{sign}(\mathcal{D})/8)}{\sqrt{|\mathcal{D}|}}C(k,s)\sum_{d\in\mathbb{Z}_{>0}}d^{-k-s}T(d^{2})f,$
where
@@
and $C(k,s)$ only depends on the weight $k$ and the variable $s$. However, a factor got lost: In formula (4.8) in [St] a factor of $d^{k-2}$ must be added. Then instead of $d^{-k-s}$ one gets $d^{-2k+2-s}$. Furthermore, the factor $e(\operatorname{sign}(\mathcal{D})/8)$ should be $e(-\operatorname{sign}(\mathcal{D})/8)$, as it was pulled out of the antilinear part of the scalar product. Then Stein’s formula coincides with ours.
@@
Let $m>6$. If $\Phi_{\mathcal{D},m,k}(f)=0$, then there exists a $p\mid N$ such that
$\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{2rk-2r-hr}}\right)f=0.$
@@
is bijective.
For a prime $p$ define
$R(p)\coloneqq\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{2rk-2r-hr}}.$
For distinct primes $p$ and $q$ the operators $R(p)$ and $R(q)$ are self-adjoint and commute, which follows from Theorem 3.2. Therefore, they are simultaneously diagonalizable for all primes $p$. Hence, if
00 0 $=\Phi_{\mathcal{D}}(f)$
$=C(m,k)\frac{e(-\operatorname{sign}(\mathcal{D})/8)}{\sqrt{|\mathcal{D}|}}\left(\sum_{\begin{subarray}{c}l=1\\ (l,N)=1\end{subarray}}^{\infty}\frac{T(l^{2})}{l^{2k-2-h}}\right)\prod_{p\mid N}\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{2rk-2r-hr}}\right)f,$
then $\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{2rk-2r-hr}}\right)f=0$ for some $p\mid N$. ∎
@@
and $M\coloneqq\bigcap_{p<\infty}(M_{p}\cap(L\otimes_{\mathbb{Z}}\mathbb{Q}))$, where $U_{p}(p)$ is the lattice $p\alpha_{1}\mathbb{Z}_{p}+\alpha_{2}\mathbb{Z}_{p}$. By [Kn, Satz 21.5] we have $M\otimes_{\mathbb{Z}}\mathbb{Z}_{p}=M_{p}$ for all $p$ and for $p\mid N$ we have
@@
By Proposition 6.2 it suffices to show that $\Phi_{\mathcal{D}}$ is injective, so assume that $\Phi_{\mathcal{D}}(f)=0$ for some $f\in\mathrm{S}_{k}(\mathcal{D})$. Let $n=\prod_{p\mid N}p$ be the radical of $N$. The genus $I\!I_{m,0}(\mathcal{D})$ is non-empty (see [N, Corollary 1.10.2]). We choose some $L\in I\!I_{m,0}(\mathcal{D})$. By Lemma 6.6 there exists a lattice $M\subset L$ such that
@@
where $n\gamma=n\mu=0$, $\operatorname{q}(\gamma)=\operatorname{q}(\mu)=0\bmod 1$ and $(\gamma,\mu)=1/n\bmod 1$. We set $H\coloneqq\langle\gamma\rangle$ so that $\mathcal{D}\cong H^{\bot}/H$. Because of (6.2) we can assume that $\mathcal{D}=H^{\bot}/H$. According to Lemma 6.3 also
$\Phi_{\widetilde{\mathcal{D}}}(\uparrow_{H}(f))=0.$
Then
$\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{2rk-2r-hr}}\right)\uparrow_{H}(f)=0$
for some $p\mid N$ by Lemma 6.5. For any $\beta\in\mathcal{D}$ clearly $(n/p)\gamma+\beta$ and $(n/p)\mu+\beta$ are not in $\widetilde{\mathcal{D}}^{p}$. If $p=2$, then $(n/2)\gamma,(n/2)\mu\in\mathcal{D}_{2}$ and
$2\operatorname{q}((n/2)\mu)+((n/2)\mu,(n/2)\gamma+\beta)$ $=2\operatorname{q}((n/2)\gamma)+((n/2)\gamma,(n/2)\mu+\beta)$
$=(n/2)^{2}(\mu,\gamma)=1/2\mod 1,$
so that $(n/2)\gamma+\beta\not\in\widetilde{\mathcal{D}}^{2*}$ and $(n/2)\mu+\beta\not\in\widetilde{\mathcal{D}}^{2*}$. Since $p\cdot((n/p)\gamma+\beta)=p\beta=p\cdot((n/p)\mu+\beta)$ and $\operatorname{q}((n/p)\gamma+\beta)=\operatorname{q}(\beta)=\operatorname{q}((n/p)\mu+\beta)$, we can apply Corollary 3.5 to obtain
$\langle\uparrow_{H}(f),e^{(n/p)\gamma+\beta}\rangle=\langle\uparrow_{H}(f),e^{(n/p)\mu+\beta}\rangle.$
Then
$\langle f,e^{\beta}\rangle$ $=\langle f,\downarrow_{H}(e^{(n/p)\gamma+\beta})\rangle=\langle\uparrow_{H}(f),e^{(n/p)\gamma+\beta}\rangle$
$=\langle\uparrow_{H}(f),e^{(n/p)\mu+\beta}\rangle=\langle f,\downarrow_{H}(e^{(n/p)\mu+\beta})\rangle=0,$
where the last equation follows from the definition of $\downarrow_{H}$ since $((n/p)\mu+\beta,\gamma)=1/p\bmod 1$ and so $(n/p)\mu+\beta\not\in H^{\bot}$. Because $\beta$ was chosen arbitrarily, it follows that $f=0$. ∎
@@
Let $H\subset\mathcal{D}$ be any isotropic subgroup such that all $p$-ranks of $H^{\bot}/H$ are less than or equal to $6$. Then any lattice in $I\!I_{m,0}(H^{\bot}/H)$ locally splits a hyperbolic plane (cf. [N, Corollary 1.9.3]). By Theorem 6.7 we know that $\mathrm{S}_{k}(H^{\bot}/H)\subset\Theta_{m,k}(H^{\bot}/H)$. It was shown in [] that

### v2

@@ 4 Vector valued Eisenstein series
In this section we will study a relation between the Eisenstein series for genus $n=2$ and the Hecke operators for $n=1$. In the scalar valued case a similar result was shown in [15] and [2], called the pullback formula. We generalize the pullback formula to the vector valued case, however using a different approach. In particular, we will prove that $\partial_{h}E_{m/2}^{(2)}\left(\left(\begin{smallmatrix}z&0\\ 0&z^{\prime}\end{smallmatrix}\right)\right)$ is the sum of the kernel functions for the Hecke operators from the previous section. When $h=0$ we get, as an additional term, the product of the genus $1$ Eisenstein series in $z$ and in $z^{\prime}$ (also cf. [32]).
@@
converges normally and thus defines a modular form (cf. [35, Theorem 1]).
@@
Let $\mathcal{D}$ be a discriminant form of even signature, $l\in\mathbb{Z}$, $A\in M_{l}$ and $B\in\mathrm{SL}_{2}(\mathbb{Z})$. Then $(AB)^{\prime}=B^{\prime}A^{\prime}$ and
@@
Let $\mathcal{D}$ be a discriminant form of even signature. Let $M\in\Gamma^{(2)}$ and $\epsilon=\operatorname{sgn}(\nu(M))$ with $\operatorname{sgn}(0)=-1$. Then
@@
Then $\psi$ defines a bijection between $\Gamma_{\infty}^{(1)}\backslash\Gamma^{(1)}\times\Gamma_{\infty}^{(1)}\backslash\Gamma^{(1)}$ and $M_{0}/\{\pm 1\}$. Let $\mathcal{D}$ be a discriminant form of even signature, then we have
@@
The next result finally describes the relation between $E_{k}^{(2)}$ and the Hecke operators and is due to Stein (cf. [32, Theorem 5.3]). We include it together with its proof because we will generalize the argument to establish Theorem 4.7.
@@
where the second sum ranges over the appropriate $A$ and $B$ (which depend on $l$). It was shown in [16, section 3.1.1] that $\partial_{h}$ commutes with the slash operator of $u(A)$ and $d(B)$ on ×$\mathbb{H}\times\mathbb{H}\subset\mathbb{H}_{2}$, i.e. for a $f:\mathbb{H}_{2}\rightarrow\mathbb{C}$ we have
@@
5 The vector valued Siegel–Weil formula
In this section we will see that the genus theta series defined in section 2 is equal to the Eisenstein series defined in section 4. This is essentially an application of the Siegel–Weil formula shown by Weil in [35]. Some missing cases were later completed by Kudla and Rallis in [20]. The Siegel–Weil formula is formulated in an adelic setup that contains the classical case that we are interested in. We will state the Siegel–Weil formula and show how to deduce the equality of the genus theta series and the Eisenstein series from it. For this purpose we will first introduce this more general setup. We mostly follow [21]. Most of the calculations in this section are known, but there seems to be no reference giving the result in the generality that we require.
@@ 5 The vector valued Siegel–Weil formula
$A(R)$ =$=\left\{a(u)=\begin{pmatrix}u&0\\ 0&(u^{T})^{-1}\end{pmatrix}\bigm{|}u\in\operatorname{GL}_{n}(R)\right\},$
$N(R)$ =$=\left\{n(s)=\begin{pmatrix}I_{n}&s\\ 0&I_{n}\end{pmatrix}\bigm{|}s\in\mathrm{Sym}_{n}(R)\right\}$
for any ring $R\supset\mathbb{Z}$. Let $V$ be a quadratic space over $\mathbb{Q}$ of dimension $m=2k$ with non-degenerate symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) and $V(R)=V\otimes_{\mathbb{Q}}R$ for any ring $R\supset\mathbb{Q}$. Let $\operatorname{O}(V)$ be the corresponding orthogonal group scheme. Then $(\mathcal{G},\mathcal{H})=(\mathrm{Sp}_{2n},\operatorname{O}(V))$ is a reductive dual pair.
@@
given by mapping a basis element $e^{\underline{\mu}}$ of $\mathbb{C}[(L^{\prime}/L)^{n}]$ to $\varphi_{\underline{\mu}}=\bigotimes_{p<\infty}\varphi_{p}$, where $\varphi_{p}\in\mathscr{S}(V(\mathbb{Q}_{p})^{n})$ is the characteristic function of $\underline{\mu}+(L\otimes\mathbb{Z}_{p})^{n}$. We obtain (cf. [38, section 2.1.3])
@@
Theorem 5.4 (Siegel–Weil formula (cf. [35] and [20])).
@@
$\chi_{\infty}(\det U)^{-1}|\det U|_{\infty}^{-k}\cdot\int_{\mathcal{H}(\mathbb{Q})\backslash\mathcal{H}(\mathbb{A})}\theta(g_{Z},h;\varphi)\mathrm{d}h$
@@
()\$\mathcal{H}(\mathbb{Q})\backslash\mathcal{H}(\mathbb{A})/S\xrightarrow{\sim}G,\qquad h\mapsto h(L\otimes\hat{\mathbb{Z}})\cap V.$
@@
()\$\int_{\mathcal{H}(\mathbb{Q})\backslash\mathcal{H}(\mathbb{A})}\theta(g_{Z},h;\varphi)\mathrm{d}h=\sum_{j}\int_{\mathcal{H}(\mathbb{Q})\backslash\mathcal{H}(\mathbb{Q})h_{j}S}\theta(g_{Z},h;\varphi)\mathrm{d}h.$
Substituting $h$ for $hh_{j}$ and applying $\mathcal{H}(\mathbb{Q})\cap h_{j}Sh_{j}^{-1}=\operatorname{Aut}(L_{j})$, we obtain
()\$\int_{\mathcal{H}(\mathbb{Q})\backslash\mathcal{H}(\mathbb{Q})h_{j}S}\theta\left(g_{Z},h;\varphi\right)\mathrm{d}h$ =()\$=\int_{\mathcal{H}(\mathbb{Q})\backslash\mathcal{H}(\mathbb{Q})h_{j}Sh_{j}^{-1}}\theta\left(g_{Z},hh_{j};\varphi\right)\mathrm{d}h$
@@
Substituting $h$ for $h_{j}hh_{j}^{-1}$, the third integral becomes
@@
where in the last equation we used that the canonical homomorphism $S\rightarrow\operatorname{O}(\mathcal{D})$ is surjective (see [25, Corollary 1.9.6.]). Combining these computations we find
$\chi_{\infty}(\det U)^{-1}$ $|\det U|_{\infty}^{-k}\cdot\int_{\mathcal{H}(\mathbb{Q})\backslash\mathcal{H}(\mathbb{A})}\theta(g_{Z},h;\varphi)\mathrm{d}h$
@@
$\mathrm{vol}(S)\cdot\mu(G)$ $=\mathrm{vol}(S)\cdot\sum_{j}\frac{1}{\#\operatorname{Aut}(L_{j})}=\mathrm{vol}(\mathcal{H}(\mathbb{Q})\backslash\mathcal{H}(\mathbb{A}))=1$
@@
For details see for example [19]. We consider the finite part. Note that for $v\in\mathbb{C}[\mathcal{D}^{n}]$, $\underline{\beta}\in\mathcal{D}^{n}$ and $\underline{x}\in\underline{\beta}+L^{n}$ we have $\iota(v)(\underline{x})=\langle v,e^{\underline{\beta}}\rangle$. Furthermore, $\omega(g_{Z})$ acts trivially on $\iota(e^{\underline{\mu}})$ since $g_{Z}\in\mathcal{G}(\mathbb{R})$ and therefore
@@
6 The space of theta series
@@
Let $\mathcal{D}$ be a discriminant form of even signature and level $N$ and let $m$ be even with $m>p\text{-rank}(\mathcal{D})$ for all primes $p$. We set $G=I\!I_{m,0}(\mathcal{D})$ and $k=m/2+h$ with $h\geq 0$. By [25, Corollary 1.10.2] $G$ is non-empty. We define a linear map
@@
Let $V$ be a $\mathbb{C}$-vector space with scalar product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) which is linear in the first variable and $(v_{i})_{i=1}^{n}\subset V$ an arbitrary finite family. Then $f:V\rightarrow V$ defined by
@@
and $g(x\cdot A)=f(g(x))$. From the first assertion we see that $\ker(g)=\ker(A)$ and since $A$ is self-adjoint, $\mathbb{C}^{n}=\operatorname{im}(A)\oplus\ker(A)=\operatorname{im}(A)\oplus\ker(g)$. This implies that
@@
We remark that in [32] the formula
$\int_{\Gamma\backslash\mathbb{H}}\langle f,E_{k}^{(2)}\left(\left(\begin{smallmatrix}z&0\\ 0&-\overline{z^{\prime}}\end{smallmatrix}\right),\overline{s}\right)\rangle y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}$ $=\frac{e(\operatorname{sign}(\mathcal{D})/8)}{\sqrt{|\mathcal{D}|}}C(k,s)\sum_{d\in\mathbb{Z}_{>0}}d^{-k-s}T(d^{2})f$
was derived, where
@@
and $C(k,s)$ only depends on the weight $k$ and the variable $s$. However, a factor got lost: In formula (4.8) in [32] a factor of $d^{k-2}$ must be added. Then instead of $d^{-k-s}$ one gets $d^{-2k+2-s}$. Furthermore, the factor $e(\operatorname{sign}(\mathcal{D})/8)$ should be $e(-\operatorname{sign}(\mathcal{D})/8)$, as it was pulled out of the antilinear part of the scalar product. Then Stein’s formula coincides with ours.
@@
Let $m>6$. If $\Phi_{\mathcal{D}}(f)=0$, then
$\prod_{p\mid N}\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{2rk-2r-hr}}\right)f=0.$
@@
is bijective. This proves the lemma. ∎
@@
and $M\coloneqq\bigcap_{p<\infty}(M_{p}\cap(L\otimes_{\mathbb{Z}}\mathbb{Q}))$, where $U_{p}(p)$ is the lattice $p\alpha_{1}\mathbb{Z}_{p}+\alpha_{2}\mathbb{Z}_{p}$. By [18, Satz 21.5] we have $M\otimes_{\mathbb{Z}}\mathbb{Z}_{p}=M_{p}$ for all $p$ and for $p\mid N$ we have
@@
By Proposition 6.2 it suffices to show that $\Phi_{\mathcal{D}}$ is injective, so assume that $\Phi_{\mathcal{D}}(f)=0$ for some $f\in\mathrm{S}_{k}(\mathcal{D})$. Let $n=\prod_{p\mid N}p$ be the radical of $N$. The genus $I\!I_{m,0}(\mathcal{D})$ is non-empty (see [25, Corollary 1.10.2]). We choose some $L\in I\!I_{m,0}(\mathcal{D})$. By Lemma 6.6 there exists a lattice $M\subset L$ such that
@@
where $n\gamma=n\mu=0$, $\operatorname{q}(\gamma)=\operatorname{q}(\mu)=0\bmod 1$ and $(\gamma,\mu)=1/n\bmod 1$. We set $H\coloneqq\langle\gamma\rangle$ so that $\mathcal{D}\cong H^{\bot}/H$. Because of (6.2), we can assume that $\mathcal{D}=H^{\bot}/H$. According to Lemma 6.3 also $\Phi_{\widetilde{\mathcal{D}}}(\uparrow_{H}(f))=0$. Then
$\prod_{p\mid N}\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{2rk-2r-hr}}\right)\uparrow_{H}(f)=0$
by Lemma 6.5. For any $\beta\in\mathcal{D}$ clearly $\gamma+\beta$ and $\mu+\beta$ are not in $\widetilde{\mathcal{D}}^{p}$ for $p\mid N$. If $2\mid N$, then $(n/2)\gamma,(n/2)\mu\in\mathcal{D}_{2}$ and
$2\operatorname{q}((n/2)\mu)+((n/2)\mu,\gamma+\beta)$ $=2\operatorname{q}((n/2)\gamma)+((n/2)\gamma,\mu+\beta)$
$=(n/2)(\mu,\gamma)=1/2\mod 1,$
so that $\gamma+\beta\not\in\widetilde{\mathcal{D}}^{2*}$ and $\mu+\beta\not\in\widetilde{\mathcal{D}}^{2*}$. Since $n\cdot(\gamma+\beta)=n\beta=n\cdot(\mu+\beta)$ and $\operatorname{q}(\gamma+\beta)=\operatorname{q}(\beta)=\operatorname{q}(\mu+\beta)$, we can apply Corollary 3.5 with set of primes $P=\{p\text{ prime}\mid p\mid N\}$. Then the element $v\coloneqq v_{\gamma+\beta,\mu+\beta,P}$ from Corollary 3.5 is given by
$v=\sum_{S\subset P}(-1)^{|S|}e^{\gamma_{S}^{\mu}+\beta}.$
Note that $(\gamma_{S}^{\mu}+\beta,\gamma)=\prod_{p\in S}\frac{1}{p}\bmod 1$, so that $\downarrow_{H}(v)=e^{\beta}$ by the definition of $\downarrow_{H}$ (see page 2). We obtain
$\langle f,e^{\beta}\rangle=\langle\uparrow_{H}(f),v\rangle=\langle\prod_{p\mid N}\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{2rk-2r-hr}}\right)\uparrow_{H}(f),v\rangle=0.$
Because $\beta$ was chosen arbitrarily, it follows that $f=0$. ∎
[…]
$\bigoplus_{q}q^{\epsilon_{q}n_{q}}$
in the notation of Conway and Sloane (cf. [10], chapter 15) and $|\mathcal{D}|=p^{\alpha}a$ with $(a,p)=1$. This is a property of the genus of $L$, rather than of $L$ itself.
@@
Let $H\subset\mathcal{D}$ be any isotropic subgroup such that all $p$-ranks of $H^{\bot}/H$ are less than or equal to $6$. Then any lattice in $I\!I_{m,0}(H^{\bot}/H)$ locally splits a hyperbolic plane (cf. [25, Corollary 1.9.3]). By Theorem 6.7 we know that $\mathrm{S}_{k}(H^{\bot}/H)\subset\Theta_{m,k}(H^{\bot}/H)$. It was shown in [24] that